# OmniGenome: Aligning RNA Sequences with Secondary Structures in Genomic Foundation Models

### v2

@@
Abstract: The alignment between RNA sequences and structures in foundation models (FMs) has yet to be thoroughly investigated. Existing FMs have struggled to establish sequence-structure alignment, hindering the free flow of genomic information between RNA sequences and structures. In this study, we introduce OmniGenome, an RNA FM trained to align RNA sequences with respect to secondary structures based on structure-contextualised modelling. The alignment enables free and bidirectional mappings between sequences and structures by utilising the flexible RNA modelling paradigm that supports versatile input and output modalities, i.e., sequence and/or structure as input/output. We implement RNA design and zero-shot secondary structure prediction as case studies to evaluate the Seq2Str and Str2Seq mapping capacity of OmniGenome. Results on the EternaV2 benchmark show that OmniGenome solved of puzzles, whereas existing FMs only solved up to of the puzzles due to the oversight of sequence-structure alignment. We leverage four comprehensive in-silico genome modelling benchmarks to evaluate performance across a diverse set of genome downstream tasks, where the results show that OmniGenome achieves state-of-the-art performance on RNA and DNA benchmarks, even without any training on DNA genomes.
Keywords: RNA Foundation Model, RNA Design, Secondary Structure Prediction
@@ 1 Introduction
RNA is a critical type of molecule that encodes a vast array of biological regulatory elements that orchestrate crucial aspects of plant growth, development, and adaptation to environmental stresses. To decipher the genomic code in RNA and manipulate RNA engineering and design, current research mainly uses bioinformatics in solving RNA genome-oriented challenges. Recent advancements in large-scale pre-trained foundation models (FMs) have demonstrated their unprecedented potential to back up existing genome analysis, as FMs are capable of learning and predicting the complex ‘genomic language’ [1] hidden in genome encoding processes. Existing FMs have been widely employed as basic sequence feature extractors to improve the performance of diverse genome analysis tasks, such as secondary structure prediction [2, 3, 4, 5], degradation rate prediction [6, 7], and mRNA vaccine design [8, 9]. In RNA, it is intriguing that the functionality and stability are intertwined with its complex structures in molecular biology [10]. However, the role of the structure as a second ‘genomic language’ to interact with sequences and solve various RNA downstream tasks has been largely ignored.
Sequence-Structure Alignment in FMs
We define alignment between sequences and secondary structures111In this paper, all structures are referred to as the secondary structures. as the bidirectional information flows. Current FMs have been struggling to establish an alignment between RNA nucleotide sequences and their folded structures, thus impeding bidirectional genomic information flows. There has been a deep scientific challenge to align RNA sequences with structures because it is not deterministic to predict sequences from structures and vice versa. In other words, an identical sequence may be folded into different sub-optimal structures because the folding patterns of RNA sequences depend on various in-vivo factors [11]. Further, a structure can be folded from different sequences composed of variational combinations of nucleotide bases, as the example shown in Fig. 1. The oversight of such alignment in existing FMs causes outstanding issues in understanding and leveraging RNA structures, such as mRNA design. For example, recent state-of-the-art RNA FMs, RNA-FM [12] and RNA-MSM [13], only solved out of puzzles in in-silico RNA design [14]. This is because they fail to decipher corresponding sequences based on structures to guide RNA design.
@@
RNA structure serves as a vital input in most of the RNA genome analysis tasks. To induce the ability of Str2Seq mapping in genomic FMs, we formulate a structure-contextualised RNA sequence reconstruction task, which stems from the representation of RNA secondary structures in texts composed of dots and brackets. As the diagram is shown in Fig. 2, we first concatenate sequence-structure pairs as inputs and then mask a small portion of nucleotide bases in the sequence. Then, we pre-train OmniGenome to reconstruct the masked nucleotide bases given the structure contexts. This simple but effective formulation of Str2Seq mapping realises structure input awareness in genomics FM pre-training and provides substantial compatibility for structure-contextualised tasks, which has been verified in our RNA design benchmark.
@@
On the other hand, Seq2Str mapping, such as end-to-end secondary structure prediction (SSP) [15, 16], is another critical aspect of achieving the alignment. We generalise end-to-end structure pre-training [17] to OmniGenome pre-training. This large-scale structure pre-training on diversified genomes supervises OmniGenome to perform Seq2Str mapping. The problem of structure pre-training lies in RNA structure annotation scarcity, which leads to biased structure predictions [18] and barriers the structure prediction robustness on small datasets. To conduct Seq2Str mapping, tremendous secondary structure annotations are required to avoid data bias. A feasible solution to RNA structure pre-training is leveraging the plausible structures calculated based on the minimum free energy. In this paper, we leverage the popular ViennaRNA [19] to serve our purpose, ‘computing’ the structures for millions of RNA sequences and perform structure pre-training in OmniGenome.
Evaluation and Results
To validate the effectiveness of OmniGenome, we designed four large-scale genome benchmarks with diverse genomics tasks. The first one is the RNA genomics benchmark (RGB) compiled in the study, which contains diverse challenging genomics understanding tasks that benefit from the sequence-structure alignment, such as degradation rate prediction. The second benchmark is the plant genomics benchmark (PGB) [20] which contains millions of DNA sequences to evaluate the DNA sequence understanding tasks. In particular, we want to use this benchmark to evaluate the generalisability of OmniGenome among diversified species and genomes. The overall performance of OmniGenome (up to M parameters) on both two benchmarks consistently outperforms existing genomics FMs with up to improvement, even compared with Agro-NT [20] which contains billion parameters. The last two benchmarks, available in the appendix, are the genomics benchmark (GB) [21] and genomics understanding evaluation (GUE) [22], which serve as two additional DNA benchmarks to evaluate generalisability on non-plant genome modelling.
In addition, we also conduct zero-shot Seq2Str and Str2Seq prediction experiments to verify the performance of sequence-structure alignment. As revealed in the experiments in Sections 3.3 and 3.3, OmniGenome achieves up to a macro-F1 score in zero-shot Seq2Str prediction, i.e., secondary structure prediction, outperforming fine-tuned FMs and bioinformatics methods like ViennaRNA. In terms of Str2Seq prediction performance, we evaluate the performance of OmniGenome in the in-silico RNA design task. We solved of complex puzzles of the EternaV2 benchmark [14], while state-of-the-art FMs such as RNA-MSM and RNA-FM only solved up to . Besides, OmniGenome only takes less than one hour to solve most of the puzzles, while most RNA design methods need to take up to hours to solve even a single puzzle.
@@
Open science is always the golden standard to promote this rising area of FM for genome modelling, which unfortunately lacks relevant high-quality resources such as code integrity, data availability, and pre-training pipeline. To address this gap, following the FAIR principles [23], we developed an open-source package222https://tinyurl.com/OmniGenome that includes step-by-step tutorials for FM pre-training and downstream tasks fine-tuning, to name a few. It provides ready-to-use genomics benchmarks and uses the API with only a few lines of code to streamline benchmarking purposes. We believe this will be a valuable resource to make this emerging AI for the RNA community to thrive.
@@ 2.1 RNA Tokenization for Alignment
We aim to implement a fine-grained alignment between RNA sequences and structures, where each base in the sequences reflects a structural label in ‘(’, ‘)’, ‘.’. Therefore, we propose an adapted implementation of the single nucleotide tokenization (SNT) method [1, 24] in OmniGenome, where the whole vocabulary, ‘A’, ‘T’, ‘C’, ‘G’, ‘U’, ‘N’, ‘(’, ‘)’, ‘.’, contains the nucleotide-level structural labels. We illustrate our tokenization based on an example shown in Fig. 3.
Our adapted SNT features bidirectional mappings between single nucleotide (SN) bases and structural labels required by sequence-structure alignment. Another reason for the adaption of SNT is that, in the realm of RNA genome modelling, the FM performance highly depends on the tokenization resolution [1, 24]. For example, the k-mers [25, 26] and BPE [27, 22] tokenization methods combine multiple bases into tokens and embeddings, which compromise modelling resolution and thus fail to the solution of fine-grained genomic tasks like structure prediction as well as base-level degrade rate prediction. Like other encoder-only models, e..g, BERT [27], we incorporated special tokens, e.g., ‘<mask>’, to implement masked language modelling.
@@ 2.3 Model Architecture
OmniGenome adopts the Transformer encoder architecture with bidirectional multi-head attention. We do not adopt recent architectures like Mamba [28, 29] and Hyena [1] because our experiments in Table 5 and Table 6 show that these architectures are not competent at RNA genome understanding. This low performance is probably because RNA sequences are much shorter than DNA sequences in the wild.
@@
No dropout is applied during pre-training, and we use the rotary position embeddings [30] to further enhance the model’s scalability to long RNA sequences.
We built a distributed training environment with Nvidia RTX GPUs, while its configuration is introduced in Appendix 6. The pre-training was finished in approximately 1 and 3 weeks for OmniGenome and OmniGenome, respectively.
@@ 2.4 Pre-training Database: OneKP
Recent studies [24, 22] have shown that data diversity can enhance FM performance without significantly increasing model capacity. For the OmniGenome pre-training, we collected transcriptome data from the OneKP initiative333https://sites.google.com/a/ualberta.ca/onekp/ [31], which compiles large-scale RNA raw sequence database from plant species. The raw sequences are not available for pre-training before processing and filtering.
@@ 2.5 Benchmark Suites
RGB is a collection of genome understanding tasks, as shown in Table 7. RGB contains SN-level tasks that are curated in this work or collected from published articles. The purpose of RGB is to benchmark FMs in challenging SN-level modelling tasks like the predictions of mRNA degradation rates and secondary structures. The sequence length in RGB ranges from to , which is enough for most RNA understanding tasks. These multi-species and SN-level tasks in RGB serve as the first comprehensive RNA benchmark to assess the modelling capabilities of FMs. For detailed information on each dataset, such as their sources and sizes, please refer to Appendix 9.1.
@@
PGB555https://huggingface.co/datasets/InstaDeepAI/plant-genomic-benchmark [20] ) shown in Table 9 provides a large-scale and comprehensive suite of DNA genome tasks designed to evaluate the modelling capabilities of FMs in plant biology. PGB involves types of DNA downstream subtasks, including a range of critical tasks such as promoter strength prediction and gene expression regression. There are datasets in total with millions of DNA sequences to be evaluated in PGB, and the sequence lengths are up to , which is very long for most of the genomic FMs. Since the original evaluation protocol is not publicly available, we have re-implemented the auto-benchmark for all the subtasks from PGB in our package. By integrating diverse downstream tasks, PGB aims to facilitate the development of plant genomics and robust assessment. Due to computational limitations, we randomly sample a maximum of k examples in all downstream datasets in PGB to evaluate the FM’s performance.
@@ 2.6 Str2Seq Modelling Case: RNA Design
One of the most challenging practices addressed by OmniGenome is RNA design, which has not been settled in existing FMs because of the oversight of sequence-structure alignment. In this section, we provide a case of RNA design that requires the exploitation of sequence-structure alignment. To address RNA design, we introduce a simple but effective genetic algorithm (GA)666The experiment script is available at: https://tinyurl.com/OmniGenome/tutorials/zeroshot_omnigenome_RNA_design.py to search feasible RNA candidates, which is based on the Str2Seq prediction capability of OmniGenome. The GA implementation details main steps in the genetic algorithm and workflow visualisation are available in Appendix 10 and Fig. 5, respectively. In Section 3.3, the experimental results on the Eterna V2 dataset indicate an impressive performance in RNA design compared to existing methods.
@@ 3.1 RNA Sequence Filtering
| Caduceus | SNT | M | B Tokens | Genome Reference Consortium | Human | DNA |
| OmniGenome | M |
@@ 3.2 Comparison Baselines
Apart from OmniGenome, we implement a plus variant, i.e., OmniGenome. In the context of OmniGenome, we assume the structure annotation from ViennaRNA is always available for enhancing the model based on structure-contextualised modelling. In SSP tasks, we can also use the ViennaRNA’s structure annotations as contexts to improve downstream SSP performance. Please refer to Appendix 8 for brief introductions of these FMs.
We can compare OmniGenome with the following RNA and DNA FMs shown in Table 2 as baselines to help evaluate the performance of OmniGenome. We are aware that some FMs are also developed for RNA, such as Uni-RNA [34], 5UTR-LM [35], etc. However, we cannot compare OmniGenome with them because their source codes are very hard to work with in our efforts or not publicly available.
@@ 3.3 Sequence-Structure Alignment Evaluation
we demonstrate the Str2Seq prediction capability of OmniGenome based on RNA design. We employed the Eterna777https://github.com/eternagame/eterna100-benchmarking/blob/master/eterna100_vienna2.txt [14] V2 benchmark, which consists of specified secondary structures. This task aims to design RNA sequences based on reference structures. We develop a genetic algorithm (GA) which exploits masked nucleotide modelling (a.k.a., masked language modelling) to find plausible RNA sequences that solve RNA design puzzles. The implementation details can be found in Fig. 5 in Appendix 10. In the GA, the population size is set at , with iterations, and the mutation rate for each base is . The evaluation metric is accuracy following existing works which indicates the number of puzzles solved by FMs. The experimental results are available in Table 3.
@@
We include a popular bioinformatics baseline of RNAInverse and select several recent DNA and RNA FMs which support masked language modelling. We exclude HyenaDNA in this experiment because it does not support masked nucleotide prediction. It is observed from Table 3 that RNAInverse solved of the RNA design puzzles, indicating a promising capability in RNA design. The FMs, such as 3UTRBERT and DNABERT2 fail in RNA design because they cannot handle SN-level modelling. Meanwhile, RNA-MSM, RNA-FM and SpliceBERT demonstrated trivial proficiency in RNA design, solving to puzzles. This observation suggests these FMs cannot precisely predict the bases without any Str2Seq prediction ability. With the help of Str2Seq, i.e., structure-contextualised sequence reconstruction, OmniGenome and OmniGenome significantly outperformed other FMs with and puzzles solved, respectively, underscoring the significance of Str2Seq in sequence-structure alignment. Besides, we expect an increase in performance with sufficient computational budgets and the findings provide crucial evidence of the significance of Str2Seq for RNA sequence design.

### v3

# Bridging Sequence-Structure Alignment in RNA Foundation Models

@@
Abstract
The alignment between RNA sequences and structures in foundation models (FMs) has yet to be thoroughly investigated. Existing FMs have struggled to establish sequence-structure alignment, hindering the free flow of genomic information between RNA sequences and structures. In this study, we introduce OmniGenome, an RNA FM trained to align RNA sequences with respect to secondary structures based on structure-contextualised modelling. The alignment enables free and bidirectional mappings between sequences and structures by utilising the flexible RNA modelling paradigm that supports versatile input and output modalities, i.e., sequence and/or structure as input/output. We implement RNA design and zero-shot secondary structure prediction as case studies to evaluate the Seq2Str and Str2Seq mapping capacity of OmniGenome. Results on the EternaV2 benchmark show that OmniGenome solved of puzzles, whereas existing FMs only solved up to of the puzzles due to the oversight of sequence-structure alignment. We leverage four comprehensive in-silico genome modelling benchmarks to evaluate performance across a diverse set of genome downstream tasks, where the results show that OmniGenome achieves state-of-the-art performance on RNA and DNA benchmarks, even without any training on DNA genomes.
@@ 1 Introduction
RNA is a critical type of molecule that encodes a vast array of biological regulatory elements that orchestrate crucial aspects of plant growth, development, and adaptation to environmental stresses. To decipher the genomic code in RNA and manipulate RNA engineering and design, current research mainly uses bioinformatics in solving RNA genome-oriented challenges. Recent advancements in large-scale pre-trained foundation models (FMs) have demonstrated their unprecedented potential to back up existing genome analysis, as FMs are capable of learning and predicting the complex ‘genomic language’ (Nguyen et al. 2023) hidden in genome encoding processes. Existing FMs have been widely employed as basic sequence feature extractors to improve the performance of diverse genome analysis tasks, such as secondary structure prediction (Tan et al. 2017; Danaee et al. 2018; Mathews 2019; Kalvari et al. 2021), degradation rate prediction (Yaish and Orenstein 2022; Wayment-Steele et al. 2022), and mRNA vaccine design (Corbett et al. 2020; Runge et al. 2023). In RNA, it is intriguing that the functionality and stability are intertwined with its complex structures in molecular biology (Ganser et al. 2019). However, the role of the structure as a second ‘genomic language’ to interact with sequences and solve various RNA downstream tasks has been largely ignored.
Sequence-Structure Alignment in GFMs
We define alignment between sequences and secondary structures111In this paper, all structures are referred to as the secondary structures. as the bidirectional information flows. Current FMs have been struggling to establish an alignment between RNA nucleotide sequences and their folded structures, thus impeding bidirectional genomic information flows. There has been a deep scientific challenge to align RNA sequences with structures because it is not deterministic to predict sequences from structures and vice versa. In other words, an identical sequence may be folded into different sub-optimal structures because the folding patterns of RNA sequences depend on various in-vivo factors (Tinoco Jr and Bustamante 1999). Further, a structure can be folded from different sequences composed of variational combinations of nucleotide bases, as the example shown in Figure 1. The oversight of such alignment in existing FMs causes outstanding issues in understanding and leveraging RNA structures, such as mRNA design. For example, recent state-of-the-art RNA FMs, RNA-FM (Chen et al. 2022) and RNA-MSM (Zhang et al. 2024), only solved out of puzzles in in-silico RNA design (Lee et al. 2014). This is because they fail to decipher corresponding sequences based on structures to guide RNA design.
@@
RNA structure serves as a vital input in most of the RNA genome analysis tasks. To induce the ability of Str2Seq mapping in genomic FMs, we formulate a structure-contextualised RNA sequence reconstruction task, which stems from the representation of RNA secondary structures in texts composed of dots and brackets. As the diagram is shown in Figure 2, we first concatenate sequence-structure pairs as inputs and then mask a small portion of nucleotide bases in the sequence. Then, we pre-train OmniGenome to reconstruct the masked nucleotide bases given the structure contexts. This simple but effective formulation of Str2Seq mapping realises structure input awareness in genomics FM pre-training and provides substantial compatibility for structure-contextualised tasks, which has been verified in our RNA design benchmark.
@@
On the other hand, Seq2Str mapping, such as end-to-end secondary structure prediction (SSP) (Sato, Akiyama, and Sakakibara 2021; Fu et al. 2022), is another critical aspect of achieving the alignment. We generalise end-to-end structure pre-training (Yan, Hamilton, and Blanchette 2022) to OmniGenome pre-training. This large-scale structure pre-training on diversified genomes supervises OmniGenome to perform Seq2Str mapping. The problem of structure pre-training lies in RNA structure annotation scarcity, which leads to biased structure predictions (Chen et al. 2020) and barriers the structure prediction robustness on small datasets. To conduct Seq2Str mapping, tremendous secondary structure annotations are required to avoid data bias. A feasible solution to RNA structure pre-training is leveraging the plausible structures calculated based on the minimum free energy. In this paper, we leverage the popular ViennaRNA (Lorenz et al. 2011) to serve our purpose, ‘computing’ the structures for millions of RNA sequences and perform structure pre-training in OmniGenome.
Evaluations and Results
To validate the effectiveness of OmniGenome, we designed four large-scale genome benchmarks with diverse genomics tasks. The first one is the RNA genomics benchmark (RGB) compiled in the study, which contains diverse challenging genomics understanding tasks that benefit from the sequence-structure alignment, such as degradation rate prediction. The second benchmark is the plant genomics benchmark (PGB) (Mendoza-Revilla et al. 2023) which contains millions of DNA sequences to evaluate the DNA sequence understanding tasks. In particular, we want to use this benchmark to evaluate the generalisability of OmniGenome among diversified species and genomes. The overall performance of OmniGenome (up to M parameters) on both two benchmarks consistently outperforms existing genomics FMs with up to improvement, even compared with Agro-NT (Mendoza-Revilla et al. 2023) which contains billion parameters. The last two benchmarks, available in the appendix, are the genomics benchmark (GB) (Grešová et al. 2023) and genomics understanding evaluation (GUE) (Zhou et al. 2023), which serve as two additional DNA benchmarks to evaluate generalisability on non-plant genome modelling.
In addition, we also conduct zero-shot Seq2Str and Str2Seq prediction experiments to verify the performance of sequence-structure alignment. As revealed in the experiments in Sections 3.3 and 3.3, OmniGenome achieves up to a macro-F1 score in zero-shot Seq2Str prediction, i.e., secondary structure prediction, outperforming fine-tuned FMs and bioinformatics methods like ViennaRNA. In terms of Str2Seq prediction performance, we evaluate the performance of OmniGenome in the in-silico RNA design task. We solved of complex puzzles of the EternaV2 benchmark (Lee et al. 2014), while state-of-the-art FMs such as RNA-MSM and RNA-FM only solved up to . Besides, OmniGenome only takes less than one hour to solve most of the puzzles, while most RNA design methods need to take up to hours to solve even a single puzzle.
@@
Open science is always the golden standard to promote this rising area of FM for genome modelling, which unfortunately lacks relevant high-quality resources such as code integrity, data availability, and pre-training pipeline. To address this gap, following the FAIR principles (Wilkinson et al. 2016), we developed an open-source package222https://github.com/yangheng95/OmniGenBench that includes step-by-step tutorials for FM pre-training and downstream tasks fine-tuning, to name a few. It provides ready-to-use genomics benchmarks and uses the API with only a few lines of code to streamline benchmarking purposes. We believe this will be a valuable resource to make this emerging AI for the RNA community to thrive.
@@ 2.1 RNA Tokenization for Alignment
We aim to implement a fine-grained alignment between RNA sequences and structures, where each base in the sequences reflects a structural label in ‘(’, ‘)’, ‘.’. Therefore, we propose an adapted implementation of the single nucleotide tokenization (SNT) method (Nguyen et al. 2023; Chen et al. 2023) in OmniGenome, where the whole vocabulary, ‘A’, ‘T’, ‘C’, ‘G’, ‘U’, ‘N’, ‘(’, ‘)’, ‘.’, contains the nucleotide-level structural labels. We illustrate our tokenization based on an example shown in Figure 3.
Our adapted SNT features bidirectional mappings between single nucleotide (SN) bases and structural labels required by sequence-structure alignment. Another reason for the adaption of SNT is that, in the realm of RNA genome modelling, the FM performance highly depends on the tokenization resolution (Nguyen et al. 2023; Chen et al. 2023). For example, the k-mers (Yang et al. 2023; Dalla-Torre et al. 2023) and BPE (Devlin et al. 2019; Zhou et al. 2023) tokenization methods combine multiple bases into tokens and embeddings, which compromise modelling resolution and thus fail to the solution of fine-grained genomic tasks like structure prediction as well as base-level degrade rate prediction. Like other encoder-only models, e..g, BERT (Devlin et al. 2019), we incorporated special tokens, e.g., ‘<mask>’, to implement masked language modelling.
@@ 2.3 Model Architecture
OmniGenome adopts the Transformer encoder architecture with bidirectional multi-head attention. We do not adopt recent architectures like Mamba (Gu and Dao 2023; Schiff et al. 2024) and Hyena (Nguyen et al. 2023) because our experiments in Table 4 and Table 5 show that these architectures are not competent at RNA genome understanding. This low performance is probably because RNA sequences are much shorter than DNA sequences in the wild.
@@
No dropout is applied during pre-training, and we use the rotary position embeddings (Su et al. 2024) to further enhance the model’s scalability to long RNA sequences.
We built a distributed training environment with Nvidia RTX GPUs, while its configuration is introduced in Appendix C. The pre-training was finished in approximately 1 and 3 weeks for OmniGenome and OmniGenome, respectively.
@@ 2.4 Pre-training Database: OneKP
Recent studies (Chen et al. 2023; Zhou et al. 2023) have shown that data diversity can enhance FM performance without significantly increasing model capacity. For the OmniGenome pre-training, we collected transcriptome data from the OneKP initiative333https://sites.google.com/a/ualberta.ca/onekp/ (Carpenter, Leebens-Mack, and et al. 2019), which compiles large-scale RNA raw sequence database from plant species. The raw sequences are not available for pre-training before processing and filtering.
@@ 2.5 Benchmark Suites
RGB is a collection of genome understanding tasks, as shown in Table 7. RGB contains SN-level tasks that are curated in this work or collected from published articles. The purpose of RGB is to benchmark FMs in challenging SN-level modelling tasks like the predictions of mRNA degradation rates and secondary structures. The sequence length in RGB ranges from to , which is enough for most RNA understanding tasks. These multi-species and SN-level tasks in RGB serve as the first comprehensive RNA benchmark to assess the modelling capabilities of FMs. For detailed information on each dataset, such as their sources and sizes, please refer to Appendix F.1.
@@
PGB555https://huggingface.co/datasets/InstaDeepAI/plant-genomic-benchmark (Mendoza-Revilla et al. 2023) ) shown in Table 9 provides a large-scale and comprehensive suite of DNA genome tasks designed to evaluate the modelling capabilities of FMs in plant biology. PGB involves types of DNA downstream subtasks, including a range of critical tasks such as promoter strength prediction and gene expression regression. There are datasets in total with millions of DNA sequences to be evaluated in PGB, and the sequence lengths are up to , which is very long for most of the genomic FMs. Since the original evaluation protocol is not publicly available, we have re-implemented the auto-benchmark for all the subtasks from PGB in our package. By integrating diverse downstream tasks, PGB aims to facilitate the development of plant genomics and robust assessment. Due to computational limitations, we randomly sample a maximum of k examples in all downstream datasets in PGB to evaluate the FM’s performance.
@@ 2.6 Str2Seq Modelling Case: RNA Design
One of the most challenging practices addressed by OmniGenome is RNA design, which has not been settled in existing FMs because of the oversight of sequence-structure alignment. In this section, we provide a case of RNA design that requires the exploitation of sequence-structure alignment. To address RNA design, we introduce a simple but effective genetic algorithm (GA)666The experiment script is available at: https://tinyurl.com/RNADesign to search feasible RNA candidates, which is based on the Str2Seq prediction capability of OmniGenome. The GA implementation details main steps in the genetic algorithm and workflow visualisation are available in Appendix G and Figure 5, respectively. In Section 3.3, the experimental results on the Eterna V2 dataset indicate an impressive performance in RNA design compared to existing methods.
@@ 3.2 Comparison Baselines
Apart from OmniGenome, we implement a plus variant, i.e., OmniGenome. In the context of OmniGenome, we assume the structure annotation from ViennaRNA is always available for enhancing the model based on structure-contextualised modelling. In SSP tasks, we can also use the ViennaRNA’s structure annotations as contexts to improve downstream SSP performance. Please refer to Appendix E for brief introductions of these FMs.
We can compare OmniGenome with the following RNA and DNA FMs shown in Table 6 as baselines to help evaluate the performance of OmniGenome. We are aware that some FMs are also developed for RNA, such as Uni-RNA (Wang et al. 2023), 5UTR-LM (Chu et al. 2024), etc. However, we cannot compare OmniGenome with them because their source codes are very hard to work with in our efforts or not publicly available.
@@ 3.3 Sequence-Structure Alignment Evaluation
we demonstrate the Str2Seq prediction capability of OmniGenome based on RNA design. We employed the Eterna (Lee et al. 2014) V2 benchmark, which consists of specified secondary structures. This task aims to design RNA sequences based on reference structures. We develop a genetic algorithm (GA) which exploits masked nucleotide modelling (a.k.a., masked language modelling) to find plausible RNA sequences that solve RNA design puzzles. The implementation details can be found in Figure 5 in Appendix G. In the GA, the population size is set at , with iterations, and the mutation rate for each base is . The evaluation metric is accuracy following existing works which indicates the number of puzzles solved by FMs. The experimental results are available in Table 2.
@@
We include a popular baseline of RNAInverse and select recent DNA and RNA FMs which support masked language modelling. We exclude HyenaDNA in this experiment because it does not support masked nucleotide prediction. It is observed from Table 2 that RNAInverse solved of the RNA design puzzles, indicating a promising capability in RNA design. The FMs, such as 3UTRBERT and DNABERT2 fail in RNA design because they cannot handle SN-level modelling. Meanwhile, RNA-MSM, RNA-FM and SpliceBERT demonstrated trivial proficiency in RNA design, solving to puzzles. This observation suggests these FMs cannot precisely predict the bases without any Str2Seq prediction ability. With the help of Str2Seq, i.e., structure-contextualised sequence reconstruction, OmniGenome and OmniGenome significantly outperformed other FMs with and puzzles solved, respectively, underscoring the significance of Str2Seq in sequence-structure alignment. Besides, we expect an increase in performance with sufficient computational budgets and the findings provide crucial evidence of the significance of Str2Seq for RNA sequence design.